The length of a rectangular playing field is 5ft less than twice it's width. If the perimeter of the playing field is 230ft, find the length and width of the field.
Answer by ptaylor(2198)

Let W=width
Then length(L)=2W-5
Perimeter(P)=2L+2W or
230=2(2W-5)+2W get rid of parens
230=4W-10+2W add 10 to both sides
230+10=4W+2W-10+10 collect like terms
240=6W divide both sides by 6
W=40 ft-------------------------------------width
2W-5=2*40-5=80-5=75 ft ------------------------length
CK
230=2(75)+2(40)=
230=150+80
230=230
also 75=2*40-5
75=75


Hope this helps----ptaylor